# PICT – A Differentiable, GPU-Accelerated Multi-Block PISO Solver for Simulation-Coupled Learning Tasks in Fluid Dynamics

### v1

@@ A.3.1 Finite Volume Method
[…]
which relates the divergence of a vector quantity $\mathrm{\mathbf{u}}$ in a finite volume $V$ to the flux $\mathrm{\mathbf{u}}\cdot\vec{n}$ over its surface $S$. In its discrete form this becomes a sum over the faces $\in\textbf{f}$ f of a discrete cell $P$
@@ A.3.4 Pressure Correction
$\begin{split}{\left[P_{F}\right]_{P}}&=\sum_{{\textbf{{D}}_{F}}}\left({\left[\overline{\alpha_{ij}A^{-1}}\right]_{{{D}_{f}}}}\right)N_{F}N_{{D}_{F}}/4,\\ \end{split}$ (31)
@@
Finally, to make the velocity divergence free, the gradient of the computed pressure is applied to $\mathrm{\mathbf{h}}$.
@@ A.5.1 Pressure Correction
$\begin{split}\partial A\left(\partial u^{**}\right)&=-\sum_{i}\left(\nabla p\right)_{i}(-1)A^{-2}\partial u^{**}_{i},\\ \end{split}$ (40)
@@
$\begin{split}\partial p^{*-1}_{P}\left(\partial h\right)&=\sum_{\textbf{F}\rightarrow j}\sum_{{\textbf{{D}}_{F}}\rightarrow k}-0.125N_{f}N_{{D}_{F}}\left({\left[\alpha_{jk}A^{-1}\right]_{F}}+{\left[\alpha_{jk}A^{-1}\right]_{{{D}_{F}}}}\right)\partial h_{{D}_{F}},\\ \end{split}$ (47)
@@ B.2 Lid-Driven Cavity
Figure 16: Velocity profiles for the 2D lid-driven cavity with Re $=100$ (top) and Re $=5000$ (bottom) for increasing resolutions. The left image of a pair is the u-velocity on the vertical center line, and the right is the v-velocity on the horizontal center line. The left pair uses a uniform grid, the right a grid that was refined towards all boundaries.
@@
Figure 17: Velocity profiles for the 3D lid-driven cavity with Re 1000 for increasing resolutions. The plots show the same quantities as in the 2D case, but the velocities are additionally normalized with the Reynolds number.
@@ B.3 Obstacle Flows
Figure 19: Schematic of the 2D vortex street geometry. The domain extends $16\,\mathrm{m}$ in the streamwise direction and $8\,\mathrm{m}$ in the transverse direction. The square obstacle, with a width of $1.5\,\mathrm{m}$ along the x-axis, is positioned $3\,\mathrm{m}$ downstream from the inlet. The obstacle’s vertical position ($y_{in}$) and its height ($y_{s}$) vary across cases, as detailed in table 2.
Here, we provide additional details about the 2D vortex street setup used in our evaluations. The computational domain extends $16~{}\mathrm{m}$ in the streamwise direction and $8\,\mathrm{m}$ in the transverse direction. The obstacle is positioned $3\,\mathrm{m}$ downstream from the inlet, with a width of $1.5\,\mathrm{m}$ along the x-axis. This geometry is visualized in figure 19. The obstacle height, denoted as $y_{s}$, varies along the y-axis across different cases, as detailed in table 2. The inlet velocity follows a Gaussian profile given by $u_{in}=u\frac{1}{{\sqrt{2\pi\sigma^{2}}}}e^{-\frac{y^{2}}{2\sigma^{2}}}$, where $u=1$ and $\sigma=0.4$. The viscosity is then set as $\nu=\frac{uy_{s}}{Re}$. An advective boundary condition is applied at the outlet, while the top and bottom walls are no-slip. To enhance accuracy, the grid is refined by a factor of 3 near the top, bottom, and around the obstacle. Additionally, the resolution is further increased in the blocks surrounding the bluff body with a ratio of 1.5. The final grid resolution varies with the obstacle height $y_{s}$, which takes values in $\{0.5,1.0,1.5,2.0\}\,\mathrm{m}$ Specifically, the corresponding grid resolutions are $268\times 132$, $268\times 136$, $268\times 140$, and $268\times 144$, with the increasing vertical extent of the computational domain with larger obstacle heights.
@@
[…]
@@
For the turbulent channel flow (TCF) scenario, we use a coordinate system with streamwise and spanwise directions along $x$ and $z$, while the wall normal direction is $y$. A dynamic forcing $\nu\left.\partial\overline{u}/\partial y\right|_{\text{wall}}$ is applied in streamwise direction to drive the flow. The mesh is a regular grid which is refined against the walls as described in the references. We consider the case of $\text{Re}_{\tau}=550$ [59] with a spatial resolution of $192\times 96\times 96$ and exponential refinement with a base of $1.095$. The domain has a physical size of $2\pi\delta\times 2\delta\times\pi\delta$ with $\delta=1$ being the channel half-width. For time stepping we used an adaptive time stepper that satisfies CFL$<$0.1, which was necessary to reduce numerical diffusion introduced by larger time steps. The velocity is initialized using a Reichardt profile [87] with small, divergence free perturbations and we set $\nu=\delta/\text{Re}_{\text{cl}}$ where $\text{Re}_{\text{cl}}$ is the expected centerline Reynolds number calculated via $\text{Re}_{\text{cl}}:=(\text{Re}_{\tau}/0.116)^{1/0.88}\approx 15037$.
The simulations were run for 20 ETT to ensure convergence before the statistics were accumulated over another 20 ETT, which results in ca. $11000t^{+}$. The resulting turbulence statistics can be found in figure 4. Using $u_{\tau}=\sqrt{\nu\left.\partial\overline{u}/\partial y\right|_{\text{wall}}}$, the non-dimensionalization is performed via $\mathrm{\mathbf{u}}^{+}=\mathrm{\mathbf{u}}/u_{\tau}$ and $y^{+}=yu_{\tau}/\delta$, and $t^{+}=tu_{\tau}^{2}/\nu$. The eddy turnover time is ETT$=tu_{\tau}/\delta$. The flow of the validation simulations is statistically stable and the averaged statistics are close to the spectral reference given sufficient resolution. The resulting averaged $\text{Re}_{\tau}=u_{\tau}/\nu$ is likewise close to the target with a value of 541.
@@ B.7 Aggregated Errors for TCF
$\begin{split}\Lambda_{\text{MSE}}&=\sum_{\lambda\in\Lambda}\frac{1}{\text{max}(|\hat{\lambda}|)}\frac{1}{Y}\sum_{y}\left|[\lambda]_{y}-[\overline{\hat{\lambda}}]_{y}\right|^{2}\Delta y,\\ \end{split}$ (64)
considering all statistics $\Lambda=\{U^{+},\overline{u^{\prime}u^{\prime}},\overline{v^{\prime}v^{\prime}},\overline{w^{\prime}w^{\prime}},\overline{u^{\prime}v^{\prime}}\}$ with equal weight. $\hat{\lambda}$ are the corresponding reference statistics [59], re-sampled at the y-locations ($[\overline{\hat{\lambda}}]_{y}$) of the computational mesh. The individual statistics are normalized with the maximum of the reference $1/\text{max}(|\hat{\lambda}|)$ to account for different magnitudes. Due to the refinement of the mesh towards the walls, the discrete integral over all sample points $y$ is a mean weighted with the cells’ size $\Delta y$. $Y$ is the total size of all cells under consideration. The resulting errors are presented in table 5.
@@ Appendix D Acknowledgements
Funding: This work was supported by the European Research Council (ERC-2019-COG #863850), and by the DFG Research Unit FOR 2987/1.

### v2

@@ A.3.1 Finite Volume Method
$\int_{V}(\nabla\cdot\mathrm{\mathbf{u}})dV=\oint_{S}(\mathrm{\mathbf{u}}\cdot\vec{n})dS,$ (22)
which relates the divergence of a vector quantity $\mathrm{\mathbf{u}}$ in a finite volume $V$ to the flux $\mathrm{\mathbf{u}}\cdot\vec{n}$ over its surface $S$. In its discrete form this becomes a sum over the faces $f\in\textbf{f}$ of a discrete cell $P$
@@ A.3.4 Pressure Correction
[…]
@@
Finally, to make the velocity divergence free, the gradient of the computed pressure is applied to $\mathrm{\mathbf{h}}$:
@@ A.5.1 Pressure Correction
[…]
@@
[…]
@@ B.2 Lid-Driven Cavity
(a) Uniform grid, Re =100=100
(b) Refined grid, Re =100=100
(c) Uniform grid, Re =5000=5000
(d) Refined grid, Re =5000=5000
Figure 16: Velocity profiles for the 2D lid-driven cavity with Re =100=100 and Re =5000=5000 for increasing resolutions. The left image of a subfigure is the u-velocity on the vertical center line, and the right is the v-velocity on the horizontal center line. (b) and (d) use a grid that was refined towards all boundaries. The reference is a high-resolution DNS [89].
@@
(a) Uniform grid
(b) Refined grid
Figure 17: Velocity profiles for the 3D lid-driven cavity with =1000=1000 for increasing resolutions. The plots show the same quantities as in the 2D case, but the velocities are additionally normalized with the Reynolds number. The reference is a high-resolution DNS [60].
@@ B.3 Obstacle Flows
Figure 19: Schematic of the 2D vortex street geometry. The domain extends $16$ in the streamwise direction and $8$ in the transverse direction. The square obstacle, with a width of $1.5$ along the x-axis, is positioned $3$ downstream from the inlet. The obstacle’s vertical position ($y_{in}$) and its height ($y_{s}$) vary across cases, as detailed in table 2.
Here, we provide additional details about the 2D vortex street setup used in our evaluations. All the geometry parameters have been non-dimensionalized with the characteristic length $y_{s}$. The computational domain extends $16$ in the streamwise direction and $8$ in the transverse direction. The obstacle is positioned $3$ downstream from the inlet, with a width of $1.5$ along the x-axis. This geometry is visualized in figure 19. The obstacle height, denoted as $y_{s}$, varies along the y-axis across different cases, as detailed in table 2. The inlet velocity follows a Gaussian profile given by $u_{in}=u\frac{1}{{\sqrt{2\pi\sigma^{2}}}}e^{-\frac{y^{2}}{2\sigma^{2}}}$, where $u=1$ and $\sigma=0.4$. The viscosity is then set as $\nu=\frac{uy_{s}}{Re}$. An advective boundary condition is applied at the outlet, while the top and bottom walls are no-slip. To enhance accuracy, the grid is refined by a factor of 3 near the top, bottom, and around the obstacle. Additionally, the resolution is further increased in the blocks surrounding the bluff body with a ratio of 1.5. The final grid resolution varies with the obstacle height $y_{s}$, which takes values in $\{0.5,1.0,1.5,2.0\}$ . Specifically, the corresponding grid resolutions are $268\times 132$, $268\times 136$, $268\times 140$, and $268\times 144$, with the increasing vertical extent of the computational domain with larger obstacle heights.
@@
Above, we tested our solver on the backward-facing step (BFS) flow with different Reynolds numbers. The BFS flow generally includes an inflow with parabolic velocity distribution and a sudden step on the lower side, as shown in figure 20. The geometry setup follows the work by Rouizi et al. [90] with a fixed $h=1$ for the height of the channel before the step. The length before the step is set as $5h$. The length after the step is set as $35h$. A suitable computational grid with refinement to the step, top and bottom wall has been chosen to ensure the critical regions have been sufficiently resolved. The refinement factor for the vertical direction is 2, resulting in the smallest grid y-size ($\Delta y=0.01h$) at the top and bottom wall and the horizontal line behind the step, while the maximum grid y-size is $\Delta y=0.02h$. For the horizontal direction, the refinement factor of the middle area is 20, resulting in a minimal grid x-size behind the step ($\Delta x=0.01h$), with the maximum ($\Delta x=0.2h$) close to the outlet. Refinement factors for the inlet and outlet blocks were adjusted to ensure smooth transitions between adjacent grid sizes, thereby avoiding abrupt changes in cell sizes. For the boundary conditions, the inlet parabolic velocity is defined by $U=6U_{\text{b}}\frac{y}{h}\left(1-\frac{y}{h}\right)$, where $U_{\text{b}}=1$. For the outlet, advective boundary conditions are used, and in order to avoid the outlet influencing the upstream area, a stabilizing buffer layer of $3h$ with slightly increased viscosity has been applied [28]. For the top, bottom, and step, the no-slip wall boundary condition has been applied. The simulation is run for a period of $tU_{b}/h=600$. Generally, the domain in front of the step is discretized using $64\times 32$ grids. Behind the step, a resolution of $544\times 128$ is employed, including a block with a resolution of $32\times 128$ for the buffer layers. As can be seem from figure 21, the flow reattachment lengths and velocity profiles resulting from our solver closely match the reference across all investigated Reynolds numbers. Training the models for the BFS scenario took 25h and 32h for $\text{NN}_{30}$ and $\text{NN}_{40}$, respectively.
Figure 20: Schematic of the 2D BFS geometry. The gap between step and top wall is $h=1$, and the inlet length is $l=5h$. The main channel has a length of $L=32h$, with a buffer region of $3h$ at the outlet. The total height of the domain is $H=h+s$, where $s$ is the step height offset below the inlet.
@@
For the turbulent channel flow (TCF) scenario, we use a coordinate system with streamwise and spanwise directions along $x$ and $z$, while the wall normal direction is $y$. A dynamic forcing $\nu\left.\partial\overline{u}/\partial y\right|_{\text{wall}}$ is applied in streamwise direction to drive the flow. The mesh is a regular grid which is refined against the walls as described in the reference. We consider the case of $\text{Re}_{\tau}=550$ [61] with a spatial resolution of $192\times 96\times 96$ and exponential refinement with a base of $1.095$. The domain has a physical size of $2\pi\delta\times 2\delta\times\pi\delta$ with $\delta=1$ being the channel half-width. For time stepping we used an adaptive time stepper that satisfies CFL<<0.1, which was necessary to reduce numerical diffusion introduced by larger time steps. The velocity is initialized using a Reichardt profile [91] with small, divergence free perturbations and we set $\nu=\delta/\text{Re}_{\text{cl}}$ where $\text{Re}_{\text{cl}}$ is the expected centerline Reynolds number calculated via $\text{Re}_{\text{cl}}:=(\text{Re}_{\tau}/0.116)^{1/0.88}\approx 15037$.
[…]
@@
B.6.1 Convergence
Figure 22: A convergence study simulating the TCF at increasing resolutions. The statistics have been non-dimentionalized with their respective average $u_{\tau}$, which approaches the expected with increasing resolution. The reference is from Hoyas and Jiménez [61].
We run a convergence study simulating the TCF at increasing resolutions, key statistics of which are shown in figure 22. Both the mean flow $U^{+}$ and the turbulent fluctuations $\overline{u_{i}^{\prime+}u_{i}^{\prime+}}$ converge towards the high-fidelity reference [61], while $\overline{u^{\prime+}v^{\prime+}}$ stays close in any case. The turbulent fluctuations do not seem to fully converge on the reference, which might be due to dissipation caused by the low order discretization scheme of PICT. While this might make the present implementation not suited for high-fidelity turbulence simulation, it is nevertheless a close approximation suitable for training hybrid solvers. The respective average $\text{Re}_{\tau}$ also converge towards the target of $\text{Re}_{\tau}=550$: $\text{Re}_{\tau,64\times 32\times 32}=467$, $\text{Re}_{\tau,96\times 48\times 48}=479$, $\text{Re}_{\tau,128\times 64\times 64}=502$, $\text{Re}_{\tau,192\times 96\times 96}=534$, $\text{Re}_{\tau,256\times 128\times 128}=541$.
B.6.2 Higher Order Moments
(a) Skewness CNN SGS
(b) Flatness CNN SGS
(c) Skewness no SGS
(d) Flatness no SGS
Figure 23: Skewness (3rd standardized moment) and flatness (4th standardized moment) of our learned CNN SGS and no SGS, accumulated over 20 ETT. The dash-dotted lines are a high-resolution $(256\times 128\times 128)$ reference simulated with PICT, which is close to the values reported by Kim et al. [1].
In addition to the mean and (co)variance, we show the higher order moments skewness and flatness of our learned CNN SGS in figure 23 and compare to a high resolution simulation run with PICT. The moments do not match the reference perfectly, but they nevertheless follow the same behavior qualitatively in the same value range. Further, it is visible that the learned CNN SGS, which matches the mean and (co)variance it was trained on almost perfectly, does not show the correct behavior near the wall for both skewness and flatness. Here it is outperformed by the no SGS variant, which does not match the lower order statistics as well. The discrepancy in the near-wall behavior is likely a consequence of the training setup. In fact, the corrector is optimized to minimize the error in the lower order statistics, without imposing any constraint on other physical quantities.
B.6.3 Temporal Two-Point Correlation
(a) Temporal correlation HR
(b) Temporal correlation CNN SGS
(c) Temporal correlation no SGS
Figure 24: Temporal two-point correlation over 2 ETT, HR has a resolution of $(256\times 128\times 128)$ while CNN SGS and no SGS use $(64\times 32\times 32)$. The plots are per-model as the available wall-normal slices differ due to the different discretization of high and low-resolution. The shown wall-normal positions have been chosen to be close to $y^{+}=17$ for all cases.
In figure 24 we plot the Temporal two-point correlation coefficient [92]
$R_{ij}(x,\Delta x,\tau)=\frac{\overline{u_{i}^{\prime}(x,t_{0})u_{j}^{\prime}(x+\Delta x,t_{0}+\tau)}}{\sqrt{\overline{u_{i}^{\prime 2}(x,t_{0})}}\sqrt{\overline{u_{j}^{\prime 2}(x+\Delta x,t_{0}+\tau)}}},$ (65)
with $\Delta x=0$. While all plots show a certain degree of noise, the PICT HR reference and the CNN SGS show significantly reduced spurious periodicity in the temporal correlation, which is present in the streamwise fluctuations of the no SGS case. The CNN SGS also needs longer to de-correlate (until ca. 0.5 ETT) than the no SGS version.
@@ B.7 Aggregated Errors for TCF
[…]
@@ Appendix D Acknowledgements
Funding: This work was supported by the European Research Council (ERC-2019-COG #863850 SpaTe), and by the DFG Research Unit FOR 2987/1.